# Quantum entanglement correlations in double quark PDFs

### v1

@@ I Introduction
does not provide a good approximation to the full dPDF [1, 2, 3, 7, 8]. Here, $i$ and $j$ denote parton flavors, and it has been assumed that the MPI involve a single hard scale $Q^{2}$, for simplicity. We shall first consider the double quark PDF at a low virtuality $Q_{0}^{2}$ of order a hadronic scale; henceforth $Q_{0}^{2}$ will be omitted from the arguments of the dPDFs. If the dPDF at the initial scale $Q_{0}^{2}$ lacks correlations then factorization of the dPDF is preserved by QCD evolution. In sec. III we perform a single evolution step towards higher $Q^{2}$ of the entire density matrix including its off-diagonal elements [^1] We note in passing that ref. [9] considered small-$x$ evolution, in the high gluon density limit, of the complete density matrix for soft gluons.. This provides first qualitative insight into double parton quantum correlations encoded in the density matrix at $Q^{2}>Q_{0}^{2}$.
@@ II.1 The Three Quark Fock State
The proton state $\left|P\right\rangle$ can be written as a coherent superposition of all possible Fock states multiplied by their respective Fock space amplitudes, the light-cone wavefunctions (LCwf). For moderate momentum fractions $x_{i}\sim 0.1$ and transverse momenta $|\vec{k}_{i}|\sim\Lambda_{\text{QCD}}$, the three valence quark state should dominate and we approximate the light-cone state of the proton by an effective three-quark wavefunction:
@@
$[{\rm d}x_{i}]=\prod_{i=1\cdots 3}\frac{{\rm d}x_{i}}{2x_{i}}\,\delta\left(1-\sum_{i}x_{i}\right)\ ,\ \ [{\rm d}^{2}k_{i}]=\prod_{i=1\cdots 3}\frac{{\rm d}^{2}k_{i}}{(2\pi)^{3}}\,(2\pi)^{3}\,\delta\left(\sum_{i}\vec{k}_{i}\right)\ ,$ (7)
@@
More involved approaches consider various model Hamiltonians with interactions, and the quark model LCwf are then obtained numerically. Ref. [38] has recently studied quark and gluon entanglement in the proton in this way. However, the wavefunctions and density matrices are then available only numerically. For this first exploration of quantum correlations associated with the entanglement negativity measure we consider the simple BS model above.
@@ II.2 Purging Entanglement Negativity with the PEN transformation
The PEN algorithm, introduced and detailed in ref. [39], can be used to remove the negativity of a density matrix. Namely, it takes a density matrix $\rho$ and produces a new density matrix $\rho^{\prime}$ such that $\mathcal{N}(\rho^{\prime})=0$ by replacing all negative eigenvalues of $\rho^{T_{2}}$ with zeroes [^3] The resulting $\rho^{\prime}$ is, indeed, a valid density matrix in that it has unit trace, and is hermitian and positive semi-definite.. In short, after diagonalizing $\rho^{T_{2}}$ the negative eigenvalues are replaced by 0; we then undo the unitary transformation to the eigenbasis of $\rho^{T_{2}}$ and the partial transposition, and finally rescale the matrix by $1/(1+\mathcal{N}(\rho))$ so that $\text{tr}\,\rho^{\prime}=1$. Note that $\rho^{\prime}$ is not guaranteed to be separable, except for small Hilbert spaces where the Peres-Horodecki criterion is sufficient to establish separability. However, since the negative eigenvalues of the partial transpose of $\rho$ are necessarily associated with quantum correlations, these specific correlations must be removed when the associated eigenvalue is sent to zero.
@@ II.3 Leveraging Center of Mass Variables
A set of suitable relative light-cone coordinates for three particles constrained to $x_{1}+x_{2}+x_{3}=1$ are [37]
@@
We can then construct the partial transpose $\tilde{\rho}^{T_{2}}$ and numerically calculate the negativity. For the BS light-cone wavefunction, we employed $\Delta\xi=\Delta\eta$ bins of width $\frac{1}{N}$ for $N=20,40,80,160$ and found that $\mathcal{N}(\tilde{\rho}_{\xi\eta,\xi^{\prime}\eta^{\prime}})$ converges to $\sim\!0.035$. This is because as the number $N$ of bins increases, the eigenvalue density of $\tilde{\rho}^{T_{2}}$ asymptotically approaches the form
@@
where $C_{i}$ is the multiplicity of the nonzero eigenvalue $\lambda_{i}$. There are $N_{\lambda}=(N+1)^{2}$ total eigenvalues of the partial transpose, but all but a few of them remain in a delta function peak at zero. The nonzero eigenvalues do not drift as $N$ changes, so increasing $N$ only increases the resolution for these values. One notable nonzero eigenvalue is $-0.03$ which provides most of the contribution to the negativity of $\tilde{\rho}$.
@@ II.4 Numerical Results for the Quark dPDF
with $O(z,y)=\bar{q}(z)\gamma^{+}q(y)$, and $q(z)$ the quark field operator at point $z$. (The notation $z^{-}+\vec{z}$ refers to a vector with LC minus component $z^{-}$, and transverse components $\vec{z}$ which represent the transverse spatial separation of the quarks.) In sec. III.3 below we verify that $f_{qq}(x_{1},x_{2})$ satisfies the dPDF DGLAP equation upon emission of a collinear gluon.
@@
The dPDF momentum sum rule [2] is satisfied [^6] That is, $\int{\rm d}x_{2}\,x_{2}f_{qq}(x_{1},x_{2})+\int{\rm d}x_{3}\,x_{3}f_{qq}(x_{1},x_{3})=(1-x_{1})f_{q}(x_{1})~{}.$ thanks to $f_{qq}(x_{1},x_{2})=f_{qq}(x_{1},1-x_{1}-x_{2})$. We also confirm that $13\int{\rm d}x_{1}\,x_{1}\,f_{q}(x_{1})=2\int[{\rm d}x_{i}]\,(x_{1}+x_{2}+x_{3})\,\frac{1}{4}\int[{\rm d}^{2}k_{i}]\,|\Psi_{\text{qqq}}(x_{1},x_{2},x_{3},\vec{k}_{1},\vec{k}_{2},\vec{k}_{3})|^{2}=\text{tr}\,\rho=$ 1.
@@
Fig. 2 shows $C(x_{1},x_{2})$ as a function of $x_{1}$ for various $x_{2}$. For small $x_{2}=0.14$, the GS model predicts weak correlations up to $x_{1}\approx 0.5$ which then turn into an anti-correlation as the boundary of phase space is approached. The correlation measure obtained from the BS wavefunction is close to the BA model for a large portion of phase space from $x_{1}\sim 0.05$ to $\sim 0.5$; the anti-correlation at small $x_{1}$, $C(x_{1},x_{2})<1$, weakens with increasing $x_{1}$. The BS wavefunction predicts a much stronger correlation for large $x_{1}\simeq 0.7$ but this extreme part of phase space is less relevant in practice. Removal of the quantum correlations associated with negativity (by application of PEN) does not affect the dPDF much when $x_{1}\gtrsim 0.1$ but it does lead to a strong correlation peak at small $x_{1}<0.1$. Hence, assuming that the GS, BA, and the BS LCwf models are close to the real dPDF of QCD we conclude that the LC quark model wavefunction requires the presence of subsystem quantum correlations so as not to strongly overpredict the dPDF at small $x_{2}$, and yet smaller $x_{1}<x_{2}$.
@@
[…]
@@ III Evolution to Higher Scales
So far we have considered the correlations of the dPDF at a low resolution scale $Q_{0}^{2}$ where a LC quark model may apply. Naturally, one may be interested in how Fig. 2 changes as we evolve to higher scales. The Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution of dPDFs to high $Q^{2}$ has been studied extensively in the literature, see e.g. [1, 2, 3, 4, 5, 7, 8, 10]. However, analysis of classical vs. quantum correlations using negativity and PEN requires knowledge of the entire density matrix, not just the diagonal elements, so we must evolve the entire density matrix. Hence, we require an extension of the dPDF DGLAP equations to off-diagonal density matrix elements. Here we consider only the first step of scale evolution where one of the three valence quarks splits into a quark and a gluon. We consider diagrams which exhibit a collinear divergence. In this section we will use expressions derived in sec. 3 of  [41], and in [42] where the quark wavefunction renormalization factor (virtual corrections) and the Fock space amplitude in LC gauge of the $|qqqg\rangle$ state (real emissions) have been determined.
@@ III.1 Corrections to the density matrix due to one gluon emission
Depending on where we insert our basis states, $\rho_{\alpha\alpha^{\prime}}=\langle P|\alpha\rangle\,\langle\alpha^{\prime}|P\rangle$, we have to consider either a gluon emitted and absorbed by the same quark in $\left|P\right\rangle$ or $\left\langle P\right|$, i.e. a virtual correction, or a “real emission” diagram where a gluon is exchanged by quark $i$ in $\left|P\right\rangle$ and the corresponding quark in $\left\langle P\right|$.
@@
×$\hskip 100.00015pt\times\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}\,(2-z-z^{\prime}+zz^{\prime})\,\frac{\Psi^{*}_{\text{qqq}}(x^{\prime}_{1},x^{\prime}_{2},\vec{k}^{\prime}_{1},\vec{k}^{\prime}_{2})}{2\sqrt{x_{1}^{\prime}x_{2}^{\prime}x_{3}^{\prime}}}\,\frac{\Psi_{\text{qqq}}(x_{1},x_{2},\vec{k}_{1},\vec{k}_{2})}{2\sqrt{x_{1}x_{2}x_{3}}}$ (30)
@@
×$\hskip 50.00008pt\times\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}\,(2-z-z^{\prime}+zz^{\prime})\,\frac{\Psi^{*}_{\text{qqq}}(x_{1}^{\prime}+x_{g},x^{\prime}_{2},\vec{k}_{1},\vec{k}_{2})}{2\sqrt{(x_{1}^{\prime}+x_{g})x_{2}^{\prime}x_{3}^{\prime}}}\,\frac{\Psi_{\text{qqq}}(x_{1}+x_{g},x_{2},\vec{k}_{1},\vec{k}_{2})}{2\sqrt{(x_{1}+x_{g})x_{2}x_{3}}}~{},$ (31)
@@
×$\hskip 85.00012pt\times(2-z-z^{\prime}+zz^{\prime})\,\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime}+x_{g},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}$ (33)
@@
×$\hskip 60.00009pt\times\left[\frac{1}{n^{2}+\Lambda^{2}}+\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime}+x_{g},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}~{}.$ (34)
@@
×$\hskip 10.00002pt\times\left[\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}-\frac{1}{2}\frac{1}{n^{2}+\Lambda^{2}}-\frac{1}{2}\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime}+x_{g},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}~{}.$ (35)
At this point we pause to verify that the trace of the density matrix is not only independent of the collinear and UV ($m_{\text{col}}$, $M_{\text{UV}}$) and soft ($x$) cutoffs but, in fact, that the ${\cal O}(\alpha_{s})$ correction cancels altogether so as to preserve the normalization $\text{tr}\,\,\rho=1$.
@@
so that $1-\eta-x_{g}=1-\tilde{\eta}$, $\eta(1-\xi)=\tilde{\eta}(1-\tilde{\xi})$, and $\xi\eta+x_{g}=\tilde{\xi}\tilde{\eta}$ (similar for the primed variables). This gives rise to a factor $\Theta(\tilde{\xi}\tilde{\eta}-x_{g})$, which allows us to extend the range of both $\tilde{\xi}$ and $\tilde{\eta}$ to $[0,1]$, and restores $z=x_{g}/\tilde{\xi}\tilde{\eta}$ as well as $\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})=\Psi_{\text{qqq}}(\tilde{\xi}\tilde{\eta},\tilde{\eta}(1-\tilde{\xi}),\vec{k}_{1},\vec{k}_{2})$. Furthermore,
@@
×$\hskip 10.00002pt\times\left[\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}-\frac{1}{2}\frac{1}{n^{2}+\Lambda^{2}}-\frac{1}{2}\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\,\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime},\eta^{\prime}(1-\xi^{\prime})+x_{g},\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta,\eta(1-\xi)+x_{g},\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}\ ,$ (39)

### v2

@@ I Introduction
does not provide a good approximation to the full dPDF [1, 2, 3, 7, 8]. Here, $i$ and $j$ denote parton flavors, and it has been assumed that the MPI involve a single hard scale $Q^{2}$, for simplicity. We shall first consider the double quark PDF at a low virtuality $Q_{0}^{2}$ of order a hadronic scale; henceforth $Q_{0}^{2}$ will be omitted from the arguments of the dPDFs. If the dPDF at the initial scale $Q_{0}^{2}$ lacks correlations, then factorization of the dPDF is preserved by QCD evolution. In sec. III we perform a single evolution step towards higher $Q^{2}$ of the entire density matrix including its off-diagonal elements [^1] We note in passing that ref. [9] considered small-$x$ evolution, in the high gluon density limit, of the complete density matrix for soft gluons.. This provides first qualitative insight into double parton quantum correlations encoded in the density matrix at $Q^{2}>Q_{0}^{2}$.
@@
The goal of this paper is to explore the application of QIT methods to reveal the potential presence of quantum correlations in double quark PDFs. In the following sec. I.1 we identify a measure for quantum correlations and a method for suppressing them in order to study the effect on the dPDF. In sec. II we introduce a simple, non-perturbative constituent quark model for the effective light-cone wave function of the proton which provides us with the entire two-quark density matrix including off-diagonal elements. Furthermore, we identify a valid “partitioning” of the reduced density matrix, in terms of unconstrained variables for the quark momentum fractions, in order to be able to study subsystem correlations with QIT methods. We discuss the convergence of the eigenspace of the discretized, partially transposed density matrix towards an infinite dimensional null space times a finite dimensional subspace for non-zero eigenvalues. Sec. III is devoted to the implementation of the first step of perturbative QCD scale evolution of the entire two-quark density matrix. Only the diagonal elements evolve in terms of convolutions of splitting functions with dPDFs but such treatment is insufficient for an analysis of the scale evolution of quantum correlations.
@@ II.1 The Three Quark Fock State
The proton state $\left|P\right\rangle$ can be written as a coherent superposition of all possible Fock states multiplied by their respective Fock space amplitudes, the light-cone wavefunctions. For moderate momentum fractions $x_{i}\sim 0.1$ and transverse momenta $|\vec{k}_{i}|\sim\Lambda_{\text{QCD}}$, the three valence quark state should dominate and we approximate the light-cone state of the proton by an effective three-quark wavefunction:
@@
${\rm d}x_{i}]=\prod_{i=1\cdots 3}\frac{{\rm d}x_{i}}{2x_{i}}\,\delta\left(1-\sum_{i}x_{i}\right)\ ,\ \ [{\rm d}^{2}k_{i}]=\prod_{i=1\cdots 3}\frac{{\rm d}^{2}k_{i}}{(2\pi)^{3}}\,(2\pi)^{3}\,\delta\left(\sum_{i}\vec{k}_{i}\right)\ ,$ (7)
@@
More involved approaches consider various model Hamiltonians with interactions, and possibly higher Fock states such as $|qqqg\rangle$ or $|qqqq\bar{q}\rangle$, and the quark model LCwf are then obtained numerically. Ref. [38] has recently studied quark and gluon entanglement in the proton in this way. However, the wavefunctions and density matrices are then available only numerically. For this first exploration of quantum correlations associated with the entanglement negativity measure we consider the simple BS model above.
@@ II.2 Purging Entanglement Negativity with the PEN transformation
The PEN (Purge Entanglement Negativity) algorithm, introduced and detailed in ref. [39], can be used to remove the negativity of a density matrix. Namely, it takes a density matrix $\rho$ and produces a new density matrix $\rho^{\prime}$ such that $\mathcal{N}(\rho^{\prime})=0$ by replacing all negative eigenvalues of $\rho^{T_{2}}$ with zeroes [^3] The resulting $\rho^{\prime}$ is, indeed, a valid density matrix in that it has unit trace, and is hermitian and positive semi-definite.. In short, after diagonalizing $\rho^{T_{2}}$ the negative eigenvalues are replaced by 0; we then undo the unitary transformation to the eigenbasis of $\rho^{T_{2}}$ and the partial transposition, and finally rescale the matrix by $1/(1+\mathcal{N}(\rho))$ so that $\text{tr}\,\rho^{\prime}=1$. Note that $\rho^{\prime}$ is not guaranteed to be separable, except for small Hilbert spaces where the Peres-Horodecki criterion is sufficient to establish separability. However, since the negative eigenvalues of the partial transpose of $\rho$ are necessarily associated with quantum correlations, these specific correlations must be removed when the associated eigenvalue is sent to zero.
@@ II.3 Leveraging Center of Mass Variables
A set of suitable relative light-cone coordinates for three particles constrained to $x_{1}+x_{2}+x_{3}=1$ is [37]
@@
We can then construct the partial transpose $\tilde{\rho}^{T_{2}}$ and numerically calculate the negativity. For the BS light-cone wavefunction, we employed $\Delta\xi=\Delta\eta$ bins of width $\frac{1}{N}$ for $N=20,40,80,160$ and found that $\mathcal{N}(\tilde{\rho}_{\xi\eta,\xi^{\prime}\eta^{\prime}})$ converges to $\sim\!0.035$. This is because once the variation of $\tilde{\rho}^{T_{2}}$ across a bin is small a further decrease of the bin width only increases the dimension of the null space of the matrix. That is, as the number $N$ of bins increases, the eigenvalue density of $\tilde{\rho}^{T_{2}}$ asymptotically approaches the form
@@
where $C_{i}$ is the multiplicity of the nonzero eigenvalue $\lambda_{i}$. This can be seen in fig. 1. There are $N_{\lambda}=(N+1)^{2}$ total eigenvalues of the partial transpose, but all but a few of them remain in a delta function peak at zero. The nonzero eigenvalues do not drift as $N$ changes, so increasing $N$ only increases the resolution for these values. One notable nonzero eigenvalue is 0.03-0.03 which provides most of the contribution to the negativity of $\tilde{\rho}$.
@@ II.4 Numerical Results for the Quark dPDF
with $O(z,y)=\bar{q}(z)\gamma^{+}q(y)$, and $q(z)$ the quark field operator at point $z$. (The notation $z^{-}+\vec{z}$ refers to a vector with LC minus component $z^{-}$, and transverse components $\vec{z}$ which represent the transverse spatial separation of the quarks.) In sec. III.3 below we verify that $f_{qq}(x_{1},x_{2})$ satisfies the dPDF Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) equation upon emission of a collinear gluon.
@@
The dPDF momentum sum rule [2] is satisfied [^6] That is, $\int{\rm d}x_{2}\,x_{2}f_{qq}(x_{1},x_{2})+\int{\rm d}x_{3}\,x_{3}f_{qq}(x_{1},x_{3})=(1-x_{1})f_{q}(x_{1})~.$ thanks to $f_{qq}(x_{1},x_{2})=f_{qq}(x_{1},1-x_{1}-x_{2})$. We also confirm that $3\int{\rm d}x_{1}\,x_{1}\,f_{q}(x_{1})=2\int[{\rm d}x_{i}]\,(x_{1}+x_{2}+x_{3})\,\frac{1}{4}\int[{\rm d}^{2}k_{i}]\,|\Psi_{\text{qqq}}(x_{1},x_{2},x_{3},\vec{k}_{1},\vec{k}_{2},\vec{k}_{3})|^{2}=\text{tr}\,\rho=1$.
Figure 2: Left: the double quark PDF of the LC quark model (see text) in the $x_{1}$-$x_{2}$ plane. Right: the ratio of the double quark PDF after removal of the entanglement negativity to the original dPDF shown on the left.
In fig. 2(left) we show the double quark PDF $f_{qq}(x_{1},x_{2})$ of the LC quark model described in sec. II.1. As expected, it peaks around $x_{1}\sim x_{2}\sim 1/3$ with tails extending to smaller $x_{i}$. The panel on the right shows the dPDF after removal of the entanglement negativity of the density matrix (via the PEN transformation), divided by the original dPDF. In much of the allowed range of $x_{1},x_{2}$ the PEN transformation has affected the dPDF by a factor of order 1, hence PEN is close to the identity map over a large portion of the domain. For $x_{1}$ or $x_{2}$ less than approximately 0.1 the dPDF has changed by a factor of 10 or more. Of course, in the region of small $x_{i}$, and for $Q^{2}$ greater than the confinement scale, one expects sea partons to dominate. In sec. III below we shall study the effects on the quark dPDF due to the perturbative emission of a gluon.
@@
Our main interest is in the nature of correlations encoded in the dPDF. Hence, in fig. 3 we plot the correlation factor $C(x_{1},x_{2})$, eq. (20), as a function of $x_{1}$ for various $x_{2}$. For small $x_{2}=0.14$, the GS model predicts weak correlations up to $x_{1}\approx 0.5$ which then turn into an anti-correlation as the boundary of phase space is approached. The correlation measure obtained from the BS wavefunction is close to the BA model for a large portion of phase space from $x_{1}\sim 0.05$ to $\sim 0.5$; the anti-correlation at small $x_{1}$, $C(x_{1},x_{2})<1$, weakens with increasing $x_{1}$. The BS wavefunction predicts a much stronger correlation for large $x_{1}\simeq 0.7$ but this extreme part of phase space is less relevant in practice. Removal of the quantum correlations associated with negativity (by application of PEN) does not affect the dPDF much when $x_{1}\gtrsim 0.1$ but it does lead to a strong correlation peak at small $x_{1}<0.1$. Hence, assuming that the GS, BA, and the BS LCwf models are close to the real dPDF of QCD we conclude that the LC quark model wavefunction requires the presence of subsystem quantum correlations so as not to strongly overpredict the dPDF at small $x_{2}$, and yet smaller $x_{1}<x_{2}$.
@@
For the panels in the first row of fig. 3 we may identify the following structure in the BS model. There exists a quantum correlated regime at small $x_{1}<x_{q}^{*}$ where the positive slope of $C(x_{1},x_{2})$ is due to quantum correlations; this regime extends up to about $x_{q}^{*}\approx 0.1$$0.15$ where the dotted and solid lines merge. Beyond $x_{q}^{*}$, correlations are classical as the dotted and solid lines are on top of each other. However, we can subdivide this further into a classically correlated regime up to $x_{c}^{*}$, which is the point where the curves reach their maximum. In the classically correlated regime between $x_{q}^{*}$ and $x_{c}^{*}$, $C(x_{1},x_{2})$ again exhibits a positive slope. Lastly, beyond $x_{c}^{*}$ we have the cutoff regime where the dPDF drops to 0 very rapidly, faster than the sPDF $f_{q}(x_{1})$, and the slope of $C(x_{1},x_{2})$ is negative. This arises due to the increasing importance of the momentum constraint.
@@ III Evolution to Higher Scales
So far we have considered the correlations of the dPDF at a low resolution scale $Q_{0}^{2}$ where a LC quark model may apply. Naturally, one may be interested in how Fig. 3 changes as we evolve to higher scales. The DGLAP evolution of dPDFs to high $Q^{2}$ has been studied extensively in the literature, see e.g. [1, 2, 3, 4, 5, 7, 8, 10]. However, analysis of classical vs. quantum correlations using negativity and PEN requires knowledge of the entire density matrix, not just the diagonal elements, so we must evolve the entire density matrix. Hence, we require an extension of the dPDF DGLAP equations to off-diagonal density matrix elements. Here we consider only the first step of scale evolution where one of the three valence quarks splits into a quark and a gluon. We consider diagrams which exhibit a collinear divergence. In this section we will use expressions derived in sec. 3 of  [41], and in [42] where the quark wavefunction renormalization factor (virtual corrections) and the Fock space amplitude in the LC gauge of the $|qqqg\rangle$ state (real emissions) have been determined.
@@ III.1 Corrections to the density matrix due to one gluon emission
Depending on where we insert our basis states, $\rho_{\alpha\alpha^{\prime}}=\langle P|\alpha\rangle\,\langle\alpha^{\prime}|P\rangle$, in fig. 4, we have to consider either a gluon emitted and absorbed by the same quark in $\left|P\right\rangle$ or $\left\langle P\right|$, i.e. a virtual correction, or a “real emission” diagram where a gluon is exchanged by quark $i$ in $\left|P\right\rangle$ and the corresponding quark in $\left\langle P\right|$.
@@
$\hskip 100.00015pt\times\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}\,(2-z-z^{\prime}+zz^{\prime})\,\frac{\Psi^{*}_{\text{qqq}}(x^{\prime}_{1},x^{\prime}_{2},\vec{k}^{\prime}_{1},\vec{k}^{\prime}_{2})}{2\sqrt{x_{1}^{\prime}x_{2}^{\prime}x_{3}^{\prime}}}\,\frac{\Psi_{\text{qqq}}(x_{1},x_{2},\vec{k}_{1},\vec{k}_{2})}{2\sqrt{x_{1}x_{2}x_{3}}}$ (30)
@@
$\hskip 50.00008pt\times\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}\,(2-z-z^{\prime}+zz^{\prime})\,\frac{\Psi^{*}_{\text{qqq}}(x_{1}^{\prime}+x_{g},x^{\prime}_{2},\vec{k}_{1},\vec{k}_{2})}{2\sqrt{(x_{1}^{\prime}+x_{g})x_{2}^{\prime}x_{3}^{\prime}}}\,\frac{\Psi_{\text{qqq}}(x_{1}+x_{g},x_{2},\vec{k}_{1},\vec{k}_{2})}{2\sqrt{(x_{1}+x_{g})x_{2}x_{3}}}~,$ (31)
@@
$\hskip 85.00012pt\times(2-z-z^{\prime}+zz^{\prime})\,\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime}+x_{g},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}$ (33)
@@
$\hskip 60.00009pt\times\left[\frac{1}{n^{2}+\Lambda^{2}}+\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime}+x_{g},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}~.$ (34)
@@
$\hskip 10.00002pt\times\left[\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}-\frac{1}{2}\frac{1}{n^{2}+\Lambda^{2}}-\frac{1}{2}\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime}+x_{g},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta+x_{g},\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}~.$ (35)
At this point we pause to verify not only that the trace of the density matrix is independent of the collinear and UV ($m_{\text{col}}$, $M_{\text{UV}}$) and soft ($x$) cutoffs but, in fact, that the ${\cal O}(\alpha_{s})$ correction cancels altogether so as to preserve the normalization $\text{tr}\,\,\rho=1$.
@@
$\frac{{\rm d}\eta\,{\rm d}\xi}{2\xi(1-\xi)\,2\eta(1-\eta)}\frac{1}{\left(1+\frac{x_{g}}{\xi\eta}\right)\left(1-\frac{x_{g}}{1-\eta}\right)}=\frac{{\rm d}\tilde{\eta}\,{\rm d}\tilde{\xi}}{2\tilde{\xi}(1-\tilde{\xi})\,2\tilde{\eta}(1-\tilde{\eta})}$ (37)
@@
$\hskip 10.00002pt\times\left[\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}-\frac{1}{2}\frac{1}{n^{2}+\Lambda^{2}}-\frac{1}{2}\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\,\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime},\eta^{\prime}(1-\xi^{\prime})+x_{g},\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta,\eta(1-\xi)+x_{g},\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}\ ,$ (39)